What is the average rate of change of the function f (x) = 9 – 5x between x = 1 and x = 2?
Answer by nerdybill(7384)

What is the average rate of change of the function f (x) = 9 – 5x between x = 1 and x = 2?
.
f(2)=9-5(2) = 9-10 = -1
f(1)=9-5(1) = 9-5 = 4
.
rate of change is:
"rise"/"run"
(f(2) - f(1))/(2-1)
(-1 - 4)/(2-1)
(-5)/(1)
-5
.